Solution to some of the exercises in previous lecture  Verification of Concurrent Software (by Jean-Jacques Lévy)

Concurrency 2

Shared Memory

Catuscia Palamidessi

INRIA Futurs and LIX - Ecole Polytechnique

The other lecturers for this course:

Jean-Jacques Lévy (INRIA Rocquencourt)

James Leifer (INRIA Rocquencourt) Eric Goubault (CEA)

http://pauillac.inria.fr/˜ leifer/teaching/mpri-concurrency-2005/

Outline

1 Solution to some of the exercises in previous lecture

Semaphores in Java Readers and Writers

Solution to some of the exercises in previous lecture  Verification of Concurrent Software (by Jean-Jacques Lévy)

Semaphores in Java

Outline

1 Solution to some of the exercises in previous lecture

Semaphores in Java Readers and Writers